supdasup:

A 10 foot long ladder is leaned against a building 8 feet up the side. How far away from the base of the building is the ladder?

supdasup:

@Pixel help

supdasup:

@JustSaiyan help

JustSaiyan:

I am not the best at Math, I'm sorry. @Vocaloid @dude @Sam

bm717:



bm717:

Does that help any bit? What do you think you do after that?

supdasup:

Honestly I dont know

bm717:

Alright. You can try to find the base by using the Pythagorean theorem. \[a ^{2}+b ^{2}=c ^{2}\]

bm717:

What numbers would go into that variable slot?

bm717:

Hint- c is the hypotenuse.

supdasup:

8^2 + 10^2 = 18^2 my best guess lol

bm717:

Actually, 10 goes at the c variable. c is always going to be the most value of a triangle, so it goes on the C.

supdasup:

I just had a lesson about this but I didnt listen, I should prob watch the recording lol

supdasup:

Oh ok thanks, that helps

bm717:

\[8^{2}+x ^{2}=10^{2}\]

bm717:

What is the x variable?

supdasup:

2?

bm717:

Lemme help you a bit more (: \[x ^{2}=10^{2}-8^{2}\]

supdasup:

wait, 2^2? cause I mean, if 10 - 8 isnt 2 then i dont know what haha

bm717:

Lol. But first we have to simplify the squared numbers.

bm717:

\[\sqrt{x}=100-64\], after you simplify \[10^{2}\] as 100 and as 64.

supdasup:

and 8^2 as 64?

bm717:

Yes lol, sorry for the mistake (:

supdasup:

oh xD thanks, I get it 36

supdasup:

so my answer is 36 or what?

bm717:

Remember, 36 must be squared. So the answer would be \[6^2\]

supdasup:

oh right lol, anyway thanks so much bm, that was a lot of help lol. Whats ur name by the way? Kinda weird calling you bm. Im Vitaly

bm717:

Lol, Sarah (:

supdasup:

Ok nice to know :))

supdasup:

Im literally gonna call u on every math question now lol

bm717:

Haha.